The accompanying data shows the demand for one type of chip used in industrial equipment from a small manufacturer.

a. Construct a chart of the data. What appears to happen when a new chip is​ introduced?

b. Develop a causal regression model to forecast demand that includes both time and the introduction of a new chip as explanatory variables.

c. What would the forecast be for the next month if a new chip is​ introduced? What would it be if a new chip is not​ introduced?

Month New Chip Introduced ooooooo Demand 3949 3544 6138 7752 10700 7766 5482 878913103 7326 8669 7219 7456 9744 14561 8564 1221419020 14751 14269 14896 15359 16527 17029
Answer #1


ANSWER:

a. chart of the data construction. What appears to happen when a new chip is​ introduced is

Regression Statistics
Multiple R 0.944266829
R Square 0.891639844
Adjusted R Square 0.881319829
Standard Error 1513.459395
Observations 24
ANOVA
df SS MS F Significance F
Regression 2 395804463.8 197902231.9 86.39908539 7.34735E-11
Residual 21 48101746.12 2290559.339
Total 23 443906210
Coefficients Standard Error t Stat P-value Lower 95% Upper 95%
Intercept 3097.183006 658.7938872 4.701292872 0.000121691 1727.146117 4467.219894
Month 535.5705134 44.68196631 11.98627898 7.43688E-11 442.6492776 628.4917491
Chip introduced 4955.863462 829.9306893 5.971418488 6.30595E-06 3229.92811 6681.798814

b. Develop a causal regression model to forecast demand that includes both time and the introduction of a new chip as explanatory variables.
so,

Demand= 3097 + 535.5*month + 4955.8*Chip introduced

c. The forecast be for the next month if a new chip is​ introduced, it would be if a new chip is not​introduced is

Forecast if new chip is introduced for month 25

= 3097 + 535.5*25 + 4955.8*1

= 21442

Forecast if new chip is not introduced for month 25

= 3097 + 535.5*25 + 4955.8*0

=16486